Operations Report – January 26th

Crew 203 Operations Report 26-JAN-2019

SOL: 14 (The final one for us here)

Name of person filing report: Freddy Castaneda

Non-nominal systems: Nothing new to report.

Notes on non-nominal systems: None

Generator (hours run): 15hr 55min; Turned on last night (25Jan2019) at 16:45; Turned off this morning (26Jan2019) at 8:40; Turned on today (26Jan2019) at 18:55.

Solar SOC – Turned on (25Jan2019) 91%; Turned off (26Jan2019) 100%; Turned on (26Jan2019) 85%

Diesel Reading – Under 1/2 tank.

Propane Reading – Over 79%

Ethanol Free Gasoline – Nothing to report.

Water (auxiliary tank) – Not in use.

Water (static tank) – About 25%; 130 gallons

Auxiliary to Static tank transfer – No

Gallons transferred: Not applicable

Water in GreenHab – About 29%; 85 gallons

Water (loft) – At level marker 12

Static to Loft Pump used – Yes; At 8:00, 17:05 and 19:10 to refill tank

Water Meter: 01405365

Toilet tank emptied: No.

Deimos rover used: No, still not functional

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: Not used.

Hours: 70.3

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Opportunity rover used: Not used.

Hours: 46.2

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Not used.

Hours: 72.2

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: Brakes on Opportunity still not working properly.

ATV’s Used: None (Honda, 300, 350.1, 350.2, 350.3)

Reason for use: None

Oil Added? No

ATV Fuel Used: None

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? Used to get crew to town for Hamburgers 🙂

CrewCar used and why, where? Used to get crew 204 from Grand Junction to MDRS

General notes and comments: Nothing to report.

Summary of internet: Nothing to report.

Summary of suits and radios: Suit #10 and Suit #6 batteries still don´t charge.

Summary of Hab operations: It was a very hard cleaning day with a delicious Hamburger as reward. Nothing new to report.

Summary of GreenHab operations: Nothing to report.

Summary of ScienceDome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns and requests to Mission Support: Nothing to report.

Our last day on Mars, it was a real pleasure for me to be here as part of crew 203, thanks for the support, I hope my work has contributed to the program. Thanks to all.

Freddy Castaneda

PS. A special greeting to my daughter, I’ll see you soon. 🙂

Greenhab Report – January 26th

Crew 203 Greenhab Report 26-Jan-2019
Greenhab Officer:
Hermes Bolivar

Environmental Control:
Heating
Cooling w/ ambient air (5 hrs)

40% Shade Cloth on
80% Shade Cloth on

Average Temperatures:
Low: 16.1°C
High: 25.4°C

Hours of Supplemental Light:
5

Daily Water Usage of Crops:
13 gallons

Water in Blue Tank (gallons):
31%, 93 Gallons.

Times of Watering for Crops:

9:37
12:30
15:40
19:15

Changes to crops:

* I use the spray for keep under control the population of gnast.

Narrative:

* In the last day the greenhab has crops of diffefent greens growing, the tomatos are be ok, the sugar peas are recover, some pots are avariable for culture. I apply the spray for reduce the population of gnast.

Result of experiment
percentage of germination of sugar peas on matian regolith

experiment (martian regolith with potting mix)
Pot 1 =20%
pot 2 =50%
Pot 3 =10%

control (potting mix only)

Pot 1 =100%
pot 2 =90%
Pot 3 =70%

Harvest:

NONE

Support/Supplies Needed:

NONE

Science Report – January 26th

Hi Mission Support this is the final science report.

1.

Title: Evaluation of germination of greens under different light wavelengths.

Research team: Hermes Bolivar, Freddy Castañeda, David Mateus

Results

We don’t have results yet, Because, the materials for build the necessary structures arrive to the station late, the experiment stay in process and David will collect the results in the next days.

2.

Topic: Star tracker Positioning systems

Researcher: Hernan David Mateus Jimenez

During the simulation it was not possible to work in this project due to time, however David is going to continue working in this project during his internship at MDRS

3.

Title: Ethnography of MDRS

Researcher: Hernan David Mateus Jimenez, Pablo Cristancho

During the simulation David sent 3 Logbooks to Pablo Cristancho and they are going to be analyzed in Colombia

4.

Topic: Recycling & Space sustainability

Researcher: Hernan David Mateus Jimenez,

During the simulation David gathered a set of data of de solid garbage produced by crew 203, this is going to be analyzed after simulation

4.

Title: Photogrammetry parameters of some samples of the MDRS region.

Researchers: Liza Forero, Fabián Saavedra.

Results: Some outcrops and rock samples were taken in situ. All the images that where obtain are being processed with other satellite images to create DEM´s (Digital Elevation Models). Each model is a 3D recreation of the photographed landscape.

5.

Title: Physic and chemical parameters of some samples of the MDRS region and sample processing with geobiologic potential.

Researchers: Liza Forero, Yael Méndez.

Results: Some grids were made in different areas of the MDRS zone, for each grid three parameters were analysed, conductivity, pH and absorbance, the results of each parameter lecture are being analysed and processed statistically and are going to be compare with an analogue in Colombia.

Samples were collected in North Pinto Hills and Beige Moon region. These samples were characterized according to their physicochemical parameters, finding that, in these places alkalophilic microorganisms can be found, with a high availability of nutrients.

6.

Title: Evaluation of microbiome from surfaces samples at the MDRS

Research team: Yael Mendez, Hermes Bolivar, Oscar Ojeda

Results:

This project could not be completed. The temperature of the incubator was not stable due to generator failures and thermal shock occurred, which affected the bacterial cultures. It is expected that next crews can resume the experiment.

7.

Title: Design and construction of an equipment for measuring, register and monitor the variables necessary for the characterization of evapotranspiration in soilless crops with simulation of regolith of Mars.

Researcher: Freddy Castañeda

Results:

An unexpected failure in the controller made the equipment unusable for taking measurements, a spare controller was requested but it didn’t reach the simulation time. The experiment will be developed on the campus of the National University of Colombia recreating the now known conditions of the Mars desert.

8.

Title: Evaluation of the germination of greens on analog Martian soil.

Researcher: Hermes Bolivar, Fredy Castañeda, David Mateus.

Results:

The project has been finished, we saw that the number of seeds on the Martian soil with the potting mix is less than the control, this result, show us the difficult for culture with the Martian soil and require of more research.

Thanks for your attention, regards,

Yael Méndez
Crew Scientist and HSO

Sol Summary Report – January 26th

Crew 203 Sol Summary Report 26-Jan-2019

Sol: 14

Summary Title: First burger on Earth, last night on Mars

Author’s name: Oscar Ojeda

Mission Status: Nominal, out of sim.

Sol Activity Summary: Our last day on the station began by a simple cereal breakfast. Then everyone took off to finish some of the experiments, and also to start organizing the station for the next crew to take over it. We had lunch, and a much-desired bath, then we headed for Hanksville for our first dinner on Earth. We can’t believe how fast time passed. We’ll probably go to bed early, leave everything as packed as possible, and get ready for our return to Earth. As of Comms window, we’re still waiting for crew 204 to arrive.

Look Ahead Plan: We’re leaving the station tomorrow morning.

Anomalies in work: None

Weather: Sunny with some clouds, low intensity to no winds.

Crew Physical Status: In good health.

EVA: None.

Reports to be filed: Sol Summary, Operations Report, Greenhab Report, Journalist Report, Science Report, HSO Report, Mission Summary.

Support Requested: None.